# Uncertainty modeling for fine-tuned implicit functions

### v1

@@ 6 Discussion and conclusion
Strengths
In this study, we advanced sparse 3D shape reconstruction for high-precision applications by introducing uncertainty modeling. We developed a flexible framework that facilitates uncertainty-aware fine-tuning of models and showcased its utility in reconstructing the lumbar spine from sparse and corrupted MRI data. Our observations suggest that traditional uncertainty methods like MC dropout might not be ideal for implicit shape reconstruction, as they tend to undermine the basic principles of implicit functions. However, our Dropsemble method effectively addresses these limitations, providing a promising alternative for uncertainty modeling.
Limitations
Despite the efficiencies Dropsembles introduce increase in resource utilization during the initial training phase, they face the same computation demands in the fine-tuning stage as Ensemble models. Another drawback is that, while using only four network samples can yield good performance, this small number introduces a moderate level of variability in results.
Future directions
In this work, we focused on uncertainty in input weights, assuming the encoder to be frozen. An interesting future direction is to combine Dropsembles with the uncertainty propagation [12] to comprehensively cover all aspects of uncertainty modeling in sparse 3D shape reconstruction.
@@ Appendix A Appendix: Elastic Weight Consolidation
Elastic Weight Consolidation (EWC) was introduced to address catastrophic forgetting in continual learning [22]. EWC is a regularization that protects crucial parameters in a network when learning a new task, in order to avoid catastrophic forgetting.
@@ Appendix B Appendix: metrics
Uncertainty estimates for regression tasks involve using unbiased estimates of the modes of the approximated posterior predictive distribution. In neural networks, this corresponds to performing multiple stochastic forward passes. For classification tasks, three main methods for uncertainty estimates are variational ratios, predictive entropy, and mutual information [68]. In our experiments, we opt for predictive entropy.
@@ Appendix C Appendix: training details
We generated two-dimensional datasets A and B with a sinusoidal decision boundary and Gaussian noise. A moderate distribution shift was modeled between the datasets by adjusting the support values of $x_{1}$: $x_{1}\in[-0.75,0.7]$ for dataset A and $x_{1}\in[-0.5,2.0]$ for dataset B. We created 1000 training samples and 500 test samples for dataset A, and 50 train and 500 test samples for dataset B. Training details could be found in Appendix C.
@@
First ReconNet encoder was trained with cross-entropy loss on a full training split of rigged atlas dataset to obtain a "frozen" encoder. We used a learning rate $0.01$ and trained for 100 epochs with early stopping.
The occupancy network architecture incorporates skip connections and ReLU activations, with dropout layers ($p=0.2$) following each linear layer except the last. To effectively model the entire lumbar spine, we adapted the strategy from [7] by replacing the learnable latent vector with an output from a pre-trained convolutional encoder. Specifically, we performed bilinear upsampling on the output of this frozen encoder to generate a detailed latent representation for each voxel. This representation, coupled with 3-dimensional voxel coordinates, was provided as input to the MLP. We used cross-entropy loss for training and fine-tuning the occupancy network. The network was trained for 100 epochs on dataset A with early stopping applied after 68 epochs. Learning rate $0.001$ and batch size 32 were used for training, where each batch we used only [64, 64, 64] random voxels. For fine-tuning on dataset B we used the same of parameters for all the methods: learning rate $0.001$ and tuned it for 50 epochs without early stopping. Results evaluated at the checkpoint of the last epoch are presented throughout the paper.
@@ Appendix E Appendix: additional experimental results
Dropsembles + EWC 92.0 $\textbf{91.5}\pm 0.1$ 9.5 6.2 19
@@
MCdropout + EWC 92.5 $91.3\pm 0.0$ 6.6 4.7 156
Dropsembles 92.6 $\textbf{92.2}\pm 0.0$ 8.2 5.3 19
Dropsembles + EWC 92.6 $\textbf{92.2}\pm 0.0$ 7.9 5.3 19

### v2

@@ 6 Discussion and conclusion
Strengths In this study, we advanced sparse 3D shape reconstruction for high-precision applications by introducing uncertainty modeling. We developed a flexible framework that facilitates uncertainty-aware fine-tuning of models and showcased its utility in reconstructing the lumbar spine from sparse and corrupted MRI data. Our observations suggest that traditional uncertainty methods like MC dropout are not ideal for implicit shape reconstruction, as they tend to undermine the basic principles of implicit functions. However, our Dropsembles method effectively addresses these limitations, providing a promising alternative for uncertainty modeling. Additionally, Dropsembles are significantly more efficient than Ensembles during the resource-heavy pre-training phase, allowing for substantial reductions in resource usage without significant loss in performance.
Limitations During the fine-tuning stage, Dropsembles face the same computation demands as Ensemble models.
Future directions In this work, we focus on the epistemic uncertainty of the decoder, assuming the encoder to be frozen. An interesting future direction is to combine Dropsembles with the uncertainty in the encoder to comprehensively cover all aspects of uncertainty modeling in sparse 3D shape reconstruction.
Acknowledgments
This study was financially supported by: 1. Personalized Health and Related Technologies (PHRT), project number 222 and 643, ETH domain, 2. ETH AI Center.
@@ Appendix A Appendix: Elastic Weight Consolidation
Elastic Weight Consolidation (EWC) was introduced to address catastrophic forgetting in continual learning [Kirkpatrick et al., 2016]. EWC is a regularization that protects crucial parameters in a network when learning a new task, in order to avoid catastrophic forgetting.
@@ Appendix B Appendix: metrics
Uncertainty estimates for regression tasks involve using unbiased estimates of the modes of the approximated posterior predictive distribution. In neural networks, this corresponds to performing multiple stochastic forward passes. For classification tasks, three main methods for uncertainty estimates are variational ratios, predictive entropy, and mutual information [Gal, 2016]. In our experiments, we opt for predictive entropy.
@@ Appendix C Appendix: training details
We generated two-dimensional datasets A and B with a sinusoidal decision boundary and Gaussian noise. A moderate distribution shift was modeled between the datasets by adjusting the support values of $x_{1}$: $x_{1}\in[-0.75,0.7]$ for dataset A and $x_{1}\in[-0.5,2.0]$ for dataset B. We created 1000 training samples and 500 test samples for dataset A, and 50 train and 500 test samples for dataset B.
@@
The convolutional encoder was trained on Dataset A and consequently frozen for all the rest of the experiments. Pooling layer was applied to produce same latent code for all 2D coordinates in a shape.
@@
ShapeNet experiment
For Dataset A, we selected 1,000 random samples from the ShapeNet dataset’s airplane category. For Dataset B, we randomly selected 10 unseen shapes and applied random erosion to simulate real-world distribution shifts. Examples of eroded and high-quality target shapes are shown in Figure 12. For both datasets, occupancies at a resolution of 128 were provided to the encoder. SDF values for Dataset A were estimated from ground-truth high-quality meshes. For Dataset B, meshes were reconstructed from the eroded shapes and subsequently converted into SDF representations.
The convolutional encoder was trained on Dataset A and consequently frozen for all the rest of the experiments. Bi-linear upsampling was applied to query latent vectors for specific 3D coordinates as proposed in convolutional occupancy networks architecture.
The 8-layer DeepSDF MLP network was trained on Dataset A for 100 epochs using a clipped $L1$ loss (with clipping parameter $\delta=0.1$), a learning rate of $0.001$, and a step scheduler as proposed in [Park et al., 2019a]. When applicable, dropout was added after each activation layer, as suggested in DeepSDF. For ensemble methods, four separate networks were trained on Dataset A. Fine-tuning on Dataset B was performed using a non-clipped $L1$ loss with a slower learning rate of $0.0001$ and limited to 75 epochs. For ensembles and Dropsembles, four network instances were employed during fine-tuning and inference. In the SIREN experiment, ReLU activations were replaced with periodic activation functions [Sitzmann et al., 2020], with dropout applied only after the 4th and 6th layers in the decoder network.
@@
First ReconNet encoder was trained with cross-entropy loss on a full training split of rigged atlas dataset to obtain a “frozen” encoder. We used a learning rate $0.01$ and trained for 100 epochs with early stopping. Bi-linear upsampling was applied to query latent vectors for specific 3D coordinates as proposed in convolutional occupancy networks architecture.
The occupancy network architecture incorporates skip connections and ReLU activations, with dropout layers ($p=0.2$) following each linear layer except the last. To effectively model the entire lumbar spine, we adapted the strategy from [Peng et al., 2020] by replacing the learnable latent vector with an output from a pretrained convolutional encoder. Specifically, we performed bilinear upsampling on the output of this frozen encoder to generate a detailed latent representation for each voxel. This representation, coupled with 3-dimensional voxel coordinates, was provided as input to the MLP. We used cross-entropy loss for training and fine-tuning the occupancy network. The network was trained for 100 epochs on dataset A with early stopping applied after 68 epochs. Learning rate $0.001$ and batch size 32 were used for training, where each batch we used only [64, 64, 64] random voxels. For fine-tuning on dataset B we used the same of parameters for all the methods: learning rate $0.001$ and tuned it for 50 epochs without early stopping. Results evaluated at the checkpoint of the last epoch are presented throughout the paper.
@@ Appendix E Appendix: additional experimental results
Figure 9: Example in the performance difference between Dropsembles w/ and w/o EWC: tiny details in patient-specific variability are not captured well in DSC metrics, but are apparent on the qualitative examples and Hausdorff distance (HD). Grey segmentations are produced by Dropsembles w/ or w/o EWC regularization overlayed with colorful segmentations of the HQ ground-truth.
Figure 10: Examples of ShapeNet samples. a) Dataset A; b) Dataset B.
Figure 11: Comparison of average predictions b/w MC Dropout and Dropsembles. Dropout predictions even after averaging are not smooth. a) Meshes produced by averaged predictions; b) A slice of SDF representation produced by averaging predicted samples.
Figure 12: Examples of Corrupted Atlas dataset.
Figure 13: Example of Atlas to MRI segmentations dataset.
Table 3: ShapeNet: Results for models trained on dataset A only. All metrics are reported in percentages [%].
Method mDICE [%] $\uparrow$ mIoU [%] $\uparrow$ ECE [%] $\downarrow$
Dropout $\underline{78.4}\pm 2.4$ $\underline{64.5}\pm 3.2$ $\textbf{31.66}\pm 3.37$
Dropout - SIREN $\textbf{78.5}\pm 3.3$ $\textbf{64.8}\pm 4.4$ $\underline{31.94}\pm 3.89$
Ensembles $76.4\pm 2.9$ $61.9\pm 3.7$ $36.33\pm 3.11$
Ensembles - SIREN $75.9\pm 3.7$ $61.3\pm 4.8$ $36.44\pm 3.85$
Table 4: ShapeNet with SIREN activations: Results for models trained on dataset A and B (fine-tuned on dataset B). All metrics are reported in percentages [%].
Method mDICE [%] $\uparrow$ mIoU [%] $\uparrow$ ECE [%] $\downarrow$
Dropout $80.2\pm 2.4$ $67.0\pm 3.3$ $27.99\pm 2.85$
Dropsembles $\underline{83.8}\pm 2.2$ $\underline{72.2}\pm 3.2$ $21.04\pm 3.04$
Ensembles $83.4\pm 2.5$ $71.6\pm 3.6$ $24.44\pm 3.03$
Dropout + EWC $83.3\pm 2.2$ $71.5\pm 3.2$ $\underline{19.27}\pm 2.74$
Dropsembles + EWC $\textbf{84.5}\pm 2.3$ $\textbf{73.2}\pm 3.4$ $\textbf{17.73}\pm 3.05$
Ensembles + EWC $80.9\pm 2.5$ $68.0\pm 3.5$ $28.95\pm 3.07$
Table 5: Detailed comparison of training time on lumbar spine experiment. We report the results in GPU-hours [h].
Method Training on atlas [h] Tuning on corrupted atlas / lumbar spine MRI [h] Total [h]
Baseline 48 - / - 48 / 48
MCdropout 48 3.5 / 22 51.5 / 99.5
Dropsembles 48 2.5 / 21 (per network) x4 58 / 132
Dropsembles + EWC 48 2.75 / 22 (per network) x4 58 / 136
Ensembles 48 (per network) x4 2.5 / 21 (per network) x4 202 / 276
@@
Dropsembles + EWC 92.0 $\textbf{91.5}\pm 0.1$ 9.5 6.1 19
Ensembles 91.9 $\underline{91.4}\pm 0.1$ 25.2 6.0 21
@@
Dropsembles 92.6 $\underline{92.2}\pm 0.0$ 8.2 5.3 19
Dropsembles + EWC 92.6 $\underline{92.2}\pm 0.0$ 6.7 5.2 19
Ensembles 92.8 $\textbf{92.5}\pm 0.1$ 6.5 5.6 21